OpenStudy (anonymous):

F -1(x) if F(x) = x + 3

OpenStudy (anonymous):

F-1(x)= x-3

OpenStudy (anonymous):

You mean : \(f^{-1}(x)\) Right??

OpenStudy (anonymous):

Let F(x) = y So: y = x + 3 Find x here: x = y - 3 Now here just flip or exchange x and y: y = x - 3 But now here : y = Inverse of F(x).. So: \(F^{-1}(x) = x - 3\).

OpenStudy (anonymous):

but it said +3

OpenStudy (anonymous):

That is \(F(x) = x + 3\) But \(F^{-1}(x) = x - 3\)

OpenStudy (anonymous):

Ok ill try again

OpenStudy (anonymous):

Go Ahead.. Keep it up..